OpenStudy (anonymous):

Determine whether the square the expression is a trinomial square. s^2-10s+25

OpenStudy (anonymous):

how u do that

OpenStudy (anonymous):

What is it about these questions that really get you confused? It could be heldful to know.

OpenStudy (immanuelv):

its binomial..... (s-5)^2

OpenStudy (anonymous):

you can tell because 25 is the square of 5 and -10 is -2*5

OpenStudy (anonymous):

for example \[x^2+12x+36\] is a perfect square because \[36=6^2\] and \[12=2\times 6\] therefore \[x^2+12+36=(x+6)^2\]